# Loss Gradient Gaussian Width based Generalization and Optimization Guarantees

### v1

@@ B.3 Population Convergence in Optimization
We provide the proof of the general mini-batch SGD algorithm, which can be applied to the GD result stated in the theorem by changing bach size $b$ to $n$.
@@
At iteration $t=(k,h)$, to simplify the notation, we use the $g(\theta_{t})=\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell\left(\theta_{(\kappa,h)},z_{j}\right)$ to be the mini-batch gradient and $\nabla{\cal L}_{{\cal D}}(\theta_{t})$ is the population gradient. Let $\Delta_{t}=g(\theta_{t})-\nabla{\cal L}_{{\cal D}}(\theta_{t})$.
@@
$\leq{\cal L}_{{\cal D}}(\theta_{t})-\frac{\eta_{t}}{2}\|\nabla{\cal L}_{{\cal D}}(\theta_{t})\|_{2}^{2}|+\frac{\eta_{t}}{2}\|\Delta_{t}\|_{2}^{2}+\tau\eta_{t}^{2}\|\nabla{\cal L}_{{\cal D}}(\theta_{t})\|+\tau\eta_{t}^{2}\|\Delta_{t}\|_{2}^{2}$
@@
From Theorem 7, with $g(\theta_{t})=\frac{1}{b}\sum_{j=(h-1)b+1}^{hb}\nabla\ell\left(\theta_{(\kappa,h)},z_{j}\right)$, and $t=(k,h)$, with probability at least $1-2\delta$, we have
$\max_{t\in T}\sup_{\theta_{t}\in\Theta}\|g(\theta_{t})-\nabla{\cal L}_{{\cal D}}(\theta_{t})\|\leq\frac{c_{1}w(\hat{\Xi}_{n}^{b})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{b}}.$ (79)
So that for any $t\in T$, with probability at least $1-2\delta$, we have
$\|\Delta_{t}\|_{2}^{2}\leq\frac{c_{1}w(\hat{\Xi}_{n}^{b})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{b}}.$ (80)
@@
where ${\cal L}_{{\cal D}}(\theta)$ is lower bounded by $L^{\star}_{{\cal D}}$. Choosing $\eta_{t}=\frac{1}{4\tau}$, we have
@@
$\leq\frac{16\tau\left({\cal L}_{{\cal D}}(\theta_{T})-L^{\star}_{{\cal D}}\right)}{T}+3\frac{c_{1}w(\hat{\Xi}_{n}^{b})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{b}}$ (83)
@@ Appendix C Gaussian Width of Gradients: Proofs for Section 4
C.1 Gaussian Width Analysis based on Loss Hessian
We focus on proving Theorem 4 and briefly discuss some extensions.
See 4
Proof.
For any $\theta\in\Theta$ and $\theta_{0}\in\mathbb{R}^{p}$, consider a quadratic approximation of the loss in terms of the loss Hessian $\hat{H}_{n}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}\ell(\theta,z_{i})$:
^$\hat{\cal L}_{n}(\theta)=\hat{\cal L}_{n}(\theta_{0})+\langle\theta-\theta_{0},\nabla\hat{\cal L}_{n}(\theta_{0})\rangle+\frac{1}{2}(\theta-\theta_{0})^{\top}\hat{H}_{n}(\tilde{\theta})(\theta-\theta_{0})$ (85)
with gradient
$\hat{\xi}_{n}(\theta)=\nabla{\cal L}_{n}(\theta)=\hat{H}_{n}(\tilde{\theta})(\theta-\theta_{0})+\nabla\hat{\cal L}_{n}(\theta_{0})~{},$ (86)
where $\tilde{\theta}=a_{\theta}\theta+(1-a_{\theta})\theta_{0}$ for some $a_{\theta}\in[0,1]$.
Consider the eigenvalue decomposition of the Hessian $\hat{H}_{n}(\tilde{\theta})=U(\tilde{\theta})\Lambda(\tilde{\theta})U(\tilde{\theta})^{\top}$. Then, the Gaussian width of gradients $\hat{\Xi}_{n}=\{\hat{\xi}_{n}(\theta),\theta\in\hat{\Theta}\}$ is given by
$w(\hat{\Xi}_{n})$ $=\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}\langle\hat{\xi}_{n}(\theta),\mathbf{g}\rangle\right]$
=$\overset{(a)}{=}\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}\langle\hat{H}_{n}(\tilde{\theta})(\theta-\theta_{0}),\mathbf{g}\rangle\right]$
$=\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}\langle U(\tilde{\theta})\Lambda(\tilde{\theta})U(\tilde{\theta})^{\top}(\theta-\theta_{0}),\mathbf{g}\rangle\right]$
$\overset{(b)}{\leq}\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}\langle\Lambda(\tilde{\theta})U(\tilde{\theta})^{\top}(\theta-\theta_{0}),\mathbf{g}\rangle\right]$
$=\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}(\theta-\theta_{0})^{\top}U(\tilde{\theta})^{\top}\Lambda(\theta)\mathbf{g}\rangle\right]$
$\overset{(c)}{\leq}\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}\|U(\tilde{\theta})(\theta-\theta_{0})\|_{2}\|\Lambda(\tilde{\theta})\mathbf{g}\|_{2}\right]$
$\overset{(d)}{\leq}c\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}\|\Lambda(\tilde{\theta})\mathbf{g}\|_{2}\right]~{},$
where (a) follows since for any set $A$ and any fixed vector $\mathbf{a}$, $w(A+\mathbf{a})=w(A)$, (b) follows since $\|U(\tilde{\theta})\|_{2}=1,\forall\tilde{\theta}$ and for any set $A$ and matrix $B$, $w(BA)\leq\|B\|_{2}w(A)$, (c) follow from Cauchy-Schwartz inequality, and (d) follows from since $\|U(\tilde{\theta})\|_{2}=1$ and $\|\theta-\theta_{0}\|_{2}\leq c$. Now, note that
$\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}\|\Lambda(\tilde{\theta})\mathbf{g}\|^{2}_{2}\right]$ $=\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}\sum_{i=1}^{p}\lambda_{i}^{2}(\tilde{\theta})g_{i}^{2}\right]$
$\overset{(a)}{\leq}\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}\sum_{i=1}^{p}\lambda_{i}^{2}(\tilde{\theta})\max_{i\in[p]}g_{i}^{2}\right]$
$=\left(\sup_{\theta}\sum_{i=1}^{p}\lambda_{i}^{2}(\tilde{\theta})\right)\mathbb{E}_{\mathbf{g}}\left[\max_{i\in[p]}g_{i}^{2}\right]$
$\overset{(b)}{\leq}\left(\sup_{\theta}\left\|\bm{\lambda}(\tilde{\theta})\right\|_{2}^{2}\right)\log p$
$\overset{(c)}{\leq}\left(\Lambda^{(u)}\right)^{2}\log p~{},$
where (a) follows by considering the worst case alignment of the ordered eigenvalues and Gaussian random variable square, (b) follows since $\mathbb{E}_{\mathbf{g}}[\max_{i\in[p]}g_{i}^{2}]=\log p$ , and (c) follows by Assumption 3.
Then, by Jensen’s inequality, we have
$\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}\|\Lambda(\tilde{\theta})\mathbf{g}\|_{2}\right]$ $=\mathbb{E}_{\mathbf{g}}\left[\left(\sup_{\theta}~{}\|\Lambda(\tilde{\theta})\mathbf{g}\|^{2}_{2}\right)^{1/2}\right]$ (87)
$\leq\left(\mathbb{E}_{\mathbf{g}}\left[\sup_{\theta}~{}\|\Lambda(\tilde{\theta})\mathbf{g}\|^{2}_{2}\right]\right)^{1/2}$ (88)
$\leq\Lambda^{(u)}\sqrt{\log p}~{}.$ (89)
That completes the proof. ∎
Based on the spectral density plots in Figure 3 from [87], it is reasonable to expect $\Lambda^{(u)}=\tilde{O}(1)$, which would imply $w(\hat{\Theta})=\tilde{O}(1)$.
Figure 3: The power-law structure of the Hessian spectrum in deep learning. The plots are from [87].
C.2 Gaussian Width bounds for Feed-Forward Networks (FFNs)
@@
$f(\theta;\mathbf{x})=\mathbf{v}^{\top}\phi(\frac{1}{\sqrt{m}}W^{(L)}\phi(\cdots\phi(\frac{1}{\sqrt{m}}W^{(1)}\mathbf{x}))))~{},$ (90)
where $W^{(1)}\in\mathbb{R}^{m\times d},W^{(l)}\in\mathbb{R}^{m\times m},l\in\{2,\ldots,L\}$ are the layer-wise weight matrices, $\mathbf{v}\in\mathbb{R}^{m}$ is the last layer vector, $\phi(\cdot)$ is the (pointwise) activation function, and the total set of parameters $\theta\in\mathbb{R}^{(L-1)m^{2}+md+m}$ is given by
@@
$\alpha^{(l)}$ $=\phi\left(\frac{1}{\sqrt{m}}W^{(l)}\alpha^{(l-1)}(\mathbf{x})\right)$ (93)
@@
As is standard, we assume $\|\mathbf{x}\|_{2}=1$ and $|y|=O(1)$.
@@
[8] Under Assumption 5, for $\theta\in B_{\rho,\rho_{1}}^{\text{Spec}}(\theta_{0})$, with probability at least $1-\frac{2}{m}$,
$\left\|W^{(l)}\right\|_{2}\leq\left(\sigma_{1}+\frac{\rho}{\sqrt{m}}\right)\sqrt{m}=\beta\sqrt{m}$
with $\beta=\sigma_{1}+\frac{\rho}{\sqrt{m}}$.
@@
Consider any $l\in\left[L\right]$. Under Assumption 4 and 5, for $\theta\in B_{\rho,\rho_{1}}^{\text{Spec}}\left(\theta_{0}\right)$, with probability at least $1-\frac{2l}{m}$, we have
$\left\|\alpha^{(l)}\right\|_{2}\leq\left(\sigma_{1}+\frac{\rho}{\sqrt{m}}\right)^{l}=\beta^{l}$
with $\beta=\sigma_{1}+\frac{\rho}{\sqrt{m}}$.
@@
$\left\|\phi\left(\frac{1}{\sqrt{m}}W^{(1)}\alpha^{(0)}\right)\right\|_{2}-\|\phi(\mathbf{0})\|_{2}\leq\left\|\phi\left(\frac{1}{\sqrt{m}}W^{(1)}\alpha^{(0)}\right)-\phi(\mathbf{0})\right\|_{2}\leq\left\|\frac{1}{\sqrt{m}}W^{(1)}\alpha^{(0)}\right\|_{2}~{},$
@@
$\|\alpha^{(1)}\|_{2}$ $=\left\|\phi\left(\frac{1}{\sqrt{m}}W^{(1)}\alpha^{(0)}\right)\right\|_{2}\leq\left\|\frac{1}{\sqrt{m}}W^{(1)}\alpha^{(0)}\right\|_{2}+\|\phi(\mathbf{0})\|_{2}$
$\leq\frac{1}{\sqrt{m}}\|W^{(1)}\|_{2}\|\alpha^{(0)}\|_{2}$
$\leq\left(\sigma_{1}+\frac{\rho}{\sqrt{m}}\right)~{},$
where we used Lemma 4 in the last inequality, which holds with probability at least $1-\frac{2}{m}$. For the inductive step, we assume that for some $l-1$, we have
$\|\alpha^{(l-1)}\|_{2}\leq\left(\sigma_{1}+\frac{\rho}{\sqrt{m}}\right)^{l-1}~{},$
which holds with the probability at least $1-\frac{2(l-1)}{m}$. Since $\phi$ is 1-Lipschitz, for layer $l$, we have
$\left\|\phi\left(\frac{1}{\sqrt{m}}W^{(l)}\alpha^{(l-1)}\right)\right\|_{2}-\|\phi(\mathbf{0})\|_{2}$ $\leq\left\|\phi\left(\frac{1}{\sqrt{m}}W^{(l)}\alpha^{(l-1)}\right)-\phi(\mathbf{0})\right\|_{2}$
$\leq\left\|\frac{1}{\sqrt{m}}W^{(l)}\alpha^{(l-1)}\right\|_{2}~{},$
@@
$\|\alpha^{(l)}\|_{2}$ $=\left\|\phi\left(\frac{1}{\sqrt{m}}W^{(l)}\alpha^{(l-1)}\right)\right\|_{2}\leq\left\|\frac{1}{\sqrt{m}}W^{(l)}\alpha^{(l-1)}\right\|_{2}+\|\phi(\mathbf{0})\|_{2}$
$\leq\frac{1}{\sqrt{m}}\|W^{(l)}\|_{2}\|\alpha^{(l-1)}\|_{2}$
$\overset{(a)}{\leq}\left(\sigma_{1}+\frac{\rho}{\sqrt{m}}\right)\|\alpha^{(l-1)}\|_{2}$
$\overset{(b)}{\leq}\left(\sigma_{1}+\frac{\rho}{\sqrt{m}}\right)^{l}~{},$
where (a) follows from Lemma 4 and (b) from the inductive step. Since we have used Lemma 4 $l$ times, after a union bound, our result would hold with probability at least $1-\frac{2l}{m}$. This completes the proof. ∎
@@
$\alpha^{(l)}=\phi\left(\tilde{\alpha}^{(l)}\right)~{},~{}~{}~{}\tilde{\alpha}^{(l)}:=\frac{1}{\sqrt{m_{l-1}}}W^{(l)}\alpha^{(l-1)}~{}.$ (96)
@@
Consider any $l\in\{2,\dots,L\}$. Under Assumptions 4 and 5, for $\theta\in B_{\rho,\rho_{1}}^{\operatorname{Spec}}(\theta_{0})$, with probability at least $\left(1-\frac{2}{m}\right)$,
$\left\|\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\right\|_{2}^{2}\leq\left(\sigma_{1}+\frac{\rho}{\sqrt{m}}\right)^{2}=\beta^{2}~{}.$ (97)
@@
$\left[\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\right]_{i,j}=\frac{1}{\sqrt{m}}\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})W_{ij}^{(l)}~{}.$ (98)
@@
$\left\|\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\right\|_{2}^{2}$ $=\sup_{\|\mathbf{u}\|_{2}=1}\frac{1}{m}\sum_{i=1}^{m}\left(\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})\sum^{m}_{j=1}W_{ij}^{(l)}u_{j}\right)^{2}$
$\overset{(a)}{\leq}\sup_{\|\mathbf{u}\|_{2}=1}\frac{1}{m}\|W^{(l)}\mathbf{u}\|_{2}^{2}$
$=\frac{1}{m}\|W^{(l)}\|_{2}^{2}$
$\overset{(b)}{\leq}\beta^{2}~{},$
@@
Consider any $l\in[L]$. Denote the parameter vector $\mathbf{w}^{(l)}=\text{vec}\left(W^{(l)}\right)$. Under Assumption 4 and 5, for $\theta\in B_{\rho,\rho_{1}}^{\text{Spec}}\left(\theta_{0}\right)$ and $\beta=\sigma_{1}+\frac{\rho}{\sqrt{m}}$, with probability at least $1-\frac{2l}{m}$,
$\left\|\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right\|_{2}^{2}\leq\frac{\beta^{2(l-1)}}{m}~{}.$

### v2

@@ B.2 Population Convergence in Optimization
At iteration $t$, to simplify the notation, we use the $g(\theta_{t})=\frac{1}{n}\sum_{i=1}^{n}\nabla\ell\left(\theta_{t},z_{i}\right)$ to be the sample gradient estimate and $\nabla{\cal L}_{{\cal D}}(\theta_{t})$ is the population gradient. Let $\Delta_{t}=g(\theta_{t})-\nabla{\cal L}_{{\cal D}}(\theta_{t})$.
@@
$\leq{\cal L}_{{\cal D}}(\theta_{t})-\frac{\eta_{t}}{2}\|\nabla{\cal L}_{{\cal D}}(\theta_{t})\|_{2}^{2}+\frac{\eta_{t}}{2}\|\Delta_{t}\|_{2}^{2}+\tau\eta_{t}^{2}\|\nabla{\cal L}_{{\cal D}}(\theta_{t})\|^{2}+\tau\eta_{t}^{2}\|\Delta_{t}\|_{2}^{2}$
@@
From Theorem 2, with probability at least $1-2\delta$, for all $t\in\{0:T\}$ we have
$\|\Delta_{t}\|_{2}^{2}\leq\frac{\left(c_{1}\max(w(\hat{\Xi}^{0}_{n}),\log p))+\sqrt{\log T+\log\frac{1}{\delta}}\right)^{2}}{n}\leq\frac{2c_{1}^{2}\max\left(w^{2}(\hat{\Xi}^{0}_{n}),\log^{2}p\right)+2\left(\log T+\log\frac{1}{\delta}\right)}{n}.$ (72)
@@
where ${\cal L}_{{\cal D}}(\theta_{T+1})$ is lower bounded by $L^{\star}_{{\cal D}}$. Choosing $\eta_{t}=\frac{1}{4\tau}$, we have
@@
$\leq\frac{16\tau\left({\cal L}_{{\cal D}}(\theta_{T})-L^{\star}_{{\cal D}}\right)}{T}+3\frac{2c_{1}^{2}\max\left(w^{2}(\hat{\Xi}^{0}_{n}),\log^{2}p\right)+2\left(\log T+\log\frac{1}{\delta}\right)}{n}$ (75)
@@ Appendix C Gaussian Width of Gradients: Proofs for Section 4
C.1 Gaussian Width bounds for Feed-Forward Networks (FFNs)
@@
$f(\theta;\mathbf{x})=\mathbf{v}^{\top}\phi(\frac{1}{\sqrt{m_{L}}}W^{(L)}\phi(\cdots\phi(\frac{1}{\sqrt{m_{1}}}W^{(1)}\mathbf{x}))))~{},$ (77)
where $W^{(1)}\in\mathbb{R}^{m_{1}\times d}$, $W^{(l)}\in\mathbb{R}^{m_{l}\times m_{l-1}},l\in\left\{2,...,L\right\}$ are layer-wise weight matrices, $\mathbf{v}\in\mathbb{R}^{m_{L}}$ is the last layer vector, $\phi(\cdot)$ is the smooth (pointwise) activation function, and the total set of parameters
$\theta:=\left(\text{vec}\left(W^{(1)}\right)^{\top},...,\text{vec}\left(W^{(L)}\right)^{\top},v^{\top}\right)^{\top}\in\mathbb{R}^{\sum_{k=1}^{L}m_{k}m_{k-1}+m_{L}}$ (78)
with $m_{0}=d$. For convenience, we write the model in terms of the layerwise outputs or features as:
@@
$\alpha^{(l)}$ $=\phi\left(\frac{1}{\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}(\mathbf{x})\right)$ (80)
@@
As is standard, we assume $\|\mathbf{x}\|_{2}=1$ according to Assumption 4.
@@
[9] Under Assumption 4, for $\theta\in B_{\rho,\rho_{1}}^{\text{Spec}}(\theta_{0})$, with probability at least $1-\frac{2}{m_{l}}$,
$\left\|W^{(l)}\right\|_{2}\leq\left(\sigma_{1}+\frac{\rho}{\sqrt{m_{l}}}\right)\sqrt{m_{l}}=\beta_{l}\sqrt{m_{l}}$
with $\beta_{l}=\sigma_{1}+\frac{\rho}{\sqrt{m_{l}}}$.
Proof.
For a ($m_{l}\times m_{l-1}$) random matrix $W_{0}^{(l)}$ with i.i.d. entries $w_{0,ij}^{(l)}\sim\mathcal{N}(0,\sigma_{0}^{(l)}{}^{2})$, with probability at least $1-2\exp(-t^{2}/2\sigma_{0}^{(l)}{}^{2})$, the largest singular value of $W_{0}$ is bounded by
$\sigma_{\max}\left(W_{0}^{(l)}\right)\leq\sqrt{m_{l}}+\sqrt{m_{l-1}}+t$
Let us choose $t=\sigma_{0}^{(l)}\sqrt{2\log m_{l}}$ so that the inequality holds with probability at least $1-\frac{2}{m_{l}}$. Then, there are two cases:
  • Case 1: $l=1$. Since $m_{0}=d$ and $m_{1}\geq d$, with probability at least $1-\frac{2}{m_{1}}$,
    $\left\|W_{0}^{(1)}\right\|_{2}\leq\sigma_{0}^{(1)}\left(\sqrt{m_{1}}+\sqrt{d}+\sqrt{2\log m_{1}}\right)\leq\sigma_{0}^{(1)}\left(2\sqrt{m_{1}}+\sqrt{2\log m_{1}}\right)=\sigma_{1}\sqrt{m_{1}}$
  • Case 2: $2\leq l\leq L$. With probability at least $1-\frac{2}{m_{l}}$,
    $\left\|W_{0}^{(l)}\right\|_{2}\leq\sigma_{0}^{(l)}\left(\sqrt{m_{l}}+\sqrt{m_{l-1}}+\sqrt{2\log m_{l}}\right)=\sigma_{1}\sqrt{m_{l}}$
Then, by triangle inequality, for $\theta\in B_{\rho,\rho_{1}}^{\text{Spec}}\left(\theta_{0}\right)$,
$\left\|W^{(l)}\right\|_{2}\leq\left\|W_{0}^{(l)}\right\|_{2}+\left\|W^{(l)}-W_{0}^{(l)}\right\|_{2}\overset{(a)}{\leq}\sigma_{1}\sqrt{m_{l}}+\rho=\beta_{l}\sqrt{m_{l}}$
@@
Consider any $l\in\left[L\right]$. Under Assumption 3 and 4, for $\theta\in B_{\rho,\rho_{1}}^{\text{Spec}}\left(\theta_{0}\right)$, with probability at least $1-\sum_{k=1}^{l}\frac{2}{m_{k}}$, we have
$\left\|\alpha^{(l)}\right\|_{2}\leq\prod_{k=1}^{l}\left(\sigma_{1}+\frac{\rho}{\sqrt{m_{k}}}\right)=\prod_{k=1}^{l}\beta_{k}$
with $\beta_{k}=\sigma_{1}+\frac{\rho}{\sqrt{m_{k}}}$.
@@
$\left\|\phi\left(\frac{1}{\sqrt{m_{1}}}W^{(1)}\alpha^{(0)}\right)\right\|_{2}-\|\phi(\mathbf{0})\|_{2}\leq\left\|\phi\left(\frac{1}{\sqrt{m_{1}}}W^{(1)}\alpha^{(0)}\right)-\phi(\mathbf{0})\right\|_{2}\leq\left\|\frac{1}{\sqrt{m_{1}}}W^{(1)}\alpha^{(0)}\right\|_{2}~{},$
@@
$\|\alpha^{(1)}\|_{2}$ $=\left\|\phi\left(\frac{1}{\sqrt{m_{1}}}W^{(1)}\alpha^{(0)}\right)\right\|_{2}\leq\left\|\frac{1}{\sqrt{m_{1}}}W^{(1)}\alpha^{(0)}\right\|_{2}+\|\phi(\mathbf{0})\|_{2}$
$\leq\frac{1}{\sqrt{m_{1}}}\|W^{(1)}\|_{2}\|\alpha^{(0)}\|_{2}$
$\leq\left(\sigma_{1}+\frac{\rho}{\sqrt{m_{1}}}\right)~{},$
where we used Lemma 6 in the last inequality, which holds with probability at least $1-\frac{2}{m_{1}}$. For the inductive step, we assume that for some $l-1$, we have
$\|\alpha^{(l-1)}\|_{2}\leq\prod_{k=1}^{l-1}\left(\sigma_{1}+\frac{\rho}{\sqrt{m_{k}}}\right)~{},$
which holds with the probability at least $1-\sum_{k=1}^{l-1}\frac{2}{m_{k}}$. Since $\phi$ is 1-Lipschitz, for layer $l$, we have
$\left\|\phi\left(\frac{1}{\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}\right)\right\|_{2}-\|\phi(\mathbf{0})\|_{2}$ $\leq\left\|\phi\left(\frac{1}{\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}\right)-\phi(\mathbf{0})\right\|_{2}$
$\leq\left\|\frac{1}{\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}\right\|_{2}~{},$
@@
$\|\alpha^{(l)}\|_{2}$ $=\left\|\phi\left(\frac{1}{\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}\right)\right\|_{2}\leq\left\|\frac{1}{\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}\right\|_{2}+\|\phi(\mathbf{0})\|_{2}$
$\leq\frac{1}{\sqrt{m_{l}}}\|W^{(l)}\|_{2}\|\alpha^{(l-1)}\|_{2}$
$\overset{(a)}{\leq}\left(\sigma_{1}+\frac{\rho}{\sqrt{m_{l}}}\right)\|\alpha^{(l-1)}\|_{2}$
$\overset{(b)}{\leq}\prod_{k=1}^{l}\left(\sigma_{1}+\frac{\rho}{\sqrt{m_{k}}}\right)~{},$
where (a) follows from Lemma 6 and (b) from the inductive step. Since we have used Lemma 6 $l$ times, after a union bound, our result would hold with probability at least $1-\sum_{k=1}^{l}\frac{2}{m_{k}}$. This completes the proof. ∎
@@
$\alpha^{(l)}=\phi\left(\tilde{\alpha}^{(l)}\right)~{},~{}~{}~{}\tilde{\alpha}^{(l)}:=\frac{1}{\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}~{}.$ (83)
@@
Consider any $l\in\{2,\dots,L\}$. Under Assumptions 3 and 4, for $\theta\in B_{\rho,\rho_{1}}^{\operatorname{Spec}}(\theta_{0})$, with probability at least $\left(1-\frac{2}{m_{l}}\right)$,
$\left\|\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\right\|_{2}^{2}\leq\left(\sigma_{1}+\frac{\rho}{\sqrt{m_{l}}}\right)^{2}=\beta_{l}^{2}~{}.$ (84)
@@
$\left[\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\right]_{i,j}=\frac{1}{\sqrt{m_{l}}}\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})W_{ij}^{(l)}~{}.$ (85)
@@
$\left\|\frac{\partial\alpha^{(l)}}{\partial\alpha^{(l-1)}}\right\|_{2}^{2}$ $=\sup_{\|\mathbf{u}\|_{2}=1}\frac{1}{m_{l}}\sum_{i=1}^{m_{l}}\left(\phi^{\prime}(\tilde{\alpha}^{(l)}_{i})\sum_{j=1}^{m_{l-1}}W_{ij}^{(l)}u_{j}\right)^{2}$
$\overset{(a)}{\leq}\sup_{\|\mathbf{u}\|_{2}=1}\frac{1}{m_{l}}\|W^{(l)}\mathbf{u}\|_{2}^{2}$
$=\frac{1}{m_{l}}\|W^{(l)}\|_{2}^{2}$
$\overset{(b)}{\leq}\beta_{l}^{2}~{},$
@@
Consider any $l\in[L]$. Denote the parameter vector $\mathbf{w}^{(l)}=\text{vec}\left(W^{(l)}\right)$. Under Assumption 3 and 4, for $\theta\in B_{\rho,\rho_{1}}^{\text{Spec}}\left(\theta_{0}\right)$ and $\beta_{k}=\sigma_{1}+\frac{\rho}{\sqrt{m_{k}}}$, with probability at least $1-\sum_{k=1}^{l-1}\frac{2}{m_{k}}$,
$\left\|\frac{\partial\alpha^{(l)}}{\partial\mathbf{w}^{(l)}}\right\|_{2}^{2}\leq\frac{\prod_{k=1}^{l-1}\beta_{k}^{2}}{m_{l}}~{}.$